differentiate log(1+x^2) with respect to tan^-1(x)
1
Expert's answer
2020-09-09T19:39:43-0400

"\frac{d}{dx}(log((1+x^2))=\frac{2x}{1+x^2}"

"\frac{d}{dx}(arctgx)=\frac{1}{1+x^2}"

Answer:

"\frac{d(log((1+x^2))}{d(arctgx)}=\frac{2x}{1+x^2}\cdot(1+x^2)=2x"